OpenStudy (anonymous):

Use a mapscale of 1/2 inch : 10 miles to find the actual distance for each measurement. 1. 5 in. ______________ 2. 2 in. ______________ 3. 7 in. ______________

OpenStudy (texaschic101):

1. ) 0.5/10 = 5/x cross multiply (.5)(x) = (10)(5) .5x = 50 x = 100 miles. 2. 0.5/10 = 2/x cross multiply (.5)(x) = (10)(2) .5x = 20 x = 40 miles 3. 0.5/10 = 7/x cross multiply (.5)(x) = (10)(7) .5x = 70 x = 140 miles

OpenStudy (texaschic101):

I set them up as proportions and I substituted .5 in for 1/2 inch.

OpenStudy (texaschic101):

Do you understand ?

OpenStudy (anonymous):

yea thanks

OpenStudy (texaschic101):

your welcome :)

OpenStudy (anonymous):

thanks